OpenStudy (anonymous):

If i have a hypotenuse number and an angle number and we are trying to find the adjacent so i know cos = adjacent / hypotenuse but since we know the angle measurements and the hypotenuse how do i write it?

OpenStudy (anonymous):

@Mathmale

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

Let the angle be v, adjacent a and the hypotenuse c. cos(v) = a/c. a = cos(v)*c. Is that what you meant?

OpenStudy (anonymous):

i already got it it was bc since my agjacent was BC so it was BC = y cos x

OpenStudy (anonymous):

@JoelSjogren but thanks anyways